solve. where appropriate, include approximations to the nearest thousandth. if no solution exists, state this
log (x + 3) + log x = 1

Answer by nerdybill(6958)

log (x + 3) + log x = 1
log x(x + 3) = 1
x(x + 3) = 10^1
x^2 + 3x = 10
x^2 + 3x - 10 = 0
(x+5)(x-2) = 0
x = {-5, 2}
Toss out the negative solution, it is an extraneous solution leaving:
x = 2